# Rethinking Efficient and Effective Point-based Networks for Event Camera Classification and Regression: EventMamba

### v2

@@ 3.4 Overall Architecture
7:
8:  Hierarchy Structure
9:  for  i in $\text{range}(m)$ do
10:     Local Extractor: Get $\mathcal{S}\in[B,T_{i-1},K,D_{i-1}]$
11:     $\mathcal{S}=LocalFE(PN)$;
12:     Temporal Aggregate: Get $\mathcal{SA}\in[B,T_{i},D_{i}]$
13:     $\mathcal{SA}=TAG(\mathcal{S})$;
14:     Global Extractor: Get $PN\in[B,T_{i},D_{i}]$
15:     $PN=GlobalFE(\mathcal{SA})$;
16:  end for
17:
18:  Classifier
19:  Get category $\hat{c}$
20:  Regressor
21:  Get 6-DOFs pose and eye coordinate
@@ 4 EXPERIMENTS
In this section, we train and test the performance of EventMamba on the server platform. The experimental evaluation encompasses assessing the model’s accuracy, mean square error (MSE) and pixel accuracy performance on the datasets, analyzing the model’s parameter size, inference time, and the number of floating-point operations (FLOPs). The platform configuration comprises the following: CPU: AMD 7950x, GPU: RTX 4090, and Memory: 32GB.
@@ 4.1 Datasets
Figure 5: Event-based action recognition, CPR and eye tracking datasets visualization. Event cameras can effectively capture moving objects, ignoring static environment.
@@ 4.2.4 Network Structure
TABLE II: Model’s Accuracy and Complexity on DVS128 Gesture.
@@ 4.3.1 DVS128 Gesture
TABLE III: Model’s Accuracy and Complexity on Daily DVS.
@@ 4.3.2 Daily DVS
TABLE IV: Model’s Accuracy and Complexity on DVS Action.
@@
TABLE V: Model’s Accuracy and Complexity on HMDB51-DVS.
@@
Network PoseNet Bayesian PoseNet Pairwise-CNN LSTM-Pose SP-LSTM CNN-LSTM PEPNet EventMamba
@@
TABLE VIII: Model’s Accuracy and Complexity on UCF101-DVS.
@@ 4.3.5 UCF101-DVS
For the larger 101-class dataset, EventMamba still performs strongly, reaching an impressive 97.1%, far exceeding all previous methods as shown in Table VIII. It is also very close to the 99.6% of the SOTA method [72] of UCF101 before synthesis. EVTC and Event-LSTM utilize the huge model and computational resources still only approach 80% accuracy. This is a great demonstration of the superiority of fine-grained temporal information and explicit temporal features.
@@
In summary, our study employed the EventMamba model to conduct an all-encompassing evaluation of action recognition datasets, ranging from small to large scales. Our analysis demonstrated that EventMamba possesses a remarkable ability to generalize and achieve outstanding performance across datasets of varying sizes. However, after capturing explicit event temporal features, the model’s representational ability is greatly enhanced and there is a greater risk of overfitting on small datasets, such as DVS Action and Daily DVS.
@@ 4.5 Results of Eye Tracking Dataset
Additionally, we conducted experiments with event-based eye tracking for the regression task. In contrast to the CPR task, which necessitates regression of translation and rotation in the world’s coordinate system, the eye tracking task solely requires regression of the pupil’s pixel coordinates within the camera coordinate system. To compare EventMamba’s performance more extensively, we use several representative Point Cloud networks [45, 10, 11, 8] for benchmarking. Table X illustrates that the optimal outcomes are highlighted in bold, whereas the second-best solutions are denoted in the underline. In comparison to classic Point Cloud networks, such as PointNet, PointNet++ and $\text{PointMLP}_{\text{elite}}$, EventMamba stands out by significantly enhancing accuracy and reducing distance error, and $\text{EventMamba}_{\text{lite}}$ also maintains high performance while consuming few computational resources. On the other hand, CNN, ConvLSTM, and CB-ConvLSTM eye-tracking frame-based methods [40] are selected for comparison. our models exhibit a remarkable improvement in tracking accuracy, with a 3% increase in the $p_{3}$ and a 6% decrease in the MSE error, while maintaining a computational cost comparable to the aforementioned methods. Besides, compared to these frame-based methods, the parameters and FLOPs in EventMamba remain constant regardless of the input camera’s resolution. That enables EventMamba to apply in different resolution edge devices.
@@ 4.6.2 Point Number
In contrast to previous work [7, 55], we increased the number of input points to 2048 on large action recognition datasets, and to determine this significant impact factor, we verified it on HMDB51-DVS. The experiments compare the accuracy at training 100 epoch as described in Table XI, EventMamba with 2048 points has 82.1% accuracy on the testset, with 1024 points has 76.6% accuracy on the testset, and with 512 points has 67.9% accuracy on the testset. This shows that increasing the number of input points can significantly increase the performance of the model on large action recognition datasets.
@@ 4.6.3 Mamba or LSTM?
Table XI and XII show the performance of the two explicit feature extractors on the HMDB51-DVS and IJRR datasets, respectively. We try to ensure that both models consume an equitable computational resource, although the Mamba-based model still slightly surpasses the LSTM-based model. On the HMDB51-DVS dataset, Mamba demonstrated a notable performance advantage, achieving an accuracy of 82.1%, which exceeded LSTM’s accuracy of 70.8% by 16%. Likewise, on the IJRR dataset, Mamba exhibited superior performance with its average Mean Squared Error (MSE) outperforming LSTM’s by 8%. Additionally, the Mamba-based model’s inference is faster than LSTM, at 5.5ms compared to 6.8ms, respectively. Ignoring the slight increase in computational resources, we believe that Mamba is superior to LSTM.
@@ 5 Conclusion
In this paper, we rethink the Point Cloud representation to address event-based tasks by introducing a spatio-temporal hierarchical structure. This innovative paradigm effectively retains fine-grained temporal information between events and aligns with the inherently sparse and asynchronous characteristics of event camera. Comprehensive experiments substantiate the efficacy of this approach in both regression and classification tasks, and we will derive this to more event-based tasks in the future.

### v4

@@ 3.4 Overall Architecture
7:  Hierarchy Structure
8:  for  i in $\text{stage}(m)$ do
9:     Local Extractor: Get $\mathcal{S}\in[B,T_{i-1},K,D_{i-1}]$
10:     $\mathcal{S}=LocalFE(PN)$;
11:     Temporal Aggregate: Get $\mathcal{SA}\in[B,T_{i},D_{i}]$
12:     $\mathcal{SA}=TAG(\mathcal{S})$;
13:     Global Extractor: Get $PN\in[B,T_{i},D_{i}]$
14:     $PN=GlobalFE(\mathcal{SA})$;
15:  end for
16:  Classifier
17:  Get category $\hat{c}$
18:  Regressor
19:  Get 6-DOFs pose and eye coordinate
@@ 4 EXPERIMENTS
In this section, we train and test the performance of EventMamba on the server platform. The experimental evaluation encompasses assessing the model’s accuracy, mean square error (MSE), and pixel accuracy performance on the datasets, analyzing the model’s parameter size, inference time, and the number of floating-point operations (FLOPs). The platform configuration comprises the following: CPU: AMD 7950x, GPU: RTX 4090, and Memory: 32GB.
@@ 4.1 Datasets
Figure 5: Event-based action recognition, CPR and eye tracking datasets visualization.
@@ 4.2.4 Network Structure
TABLE II: Accuracy and Complexity on DVS128 Gesture.
@@ 4.3.1 DVS128 Gesture
TABLE III: Accuracy and Complexity on Daily DVS.
@@ 4.3.2 Daily DVS
TABLE IV: Accuracy and Complexity on DVS Action.
@@
TABLE V: Accuracy and Complexity on HMDB51-DVS.
@@
Network PoseNet [19] Bayesian PoseNet [74] Pairwise-CNN [75] LSTM-Pose [21] SP-LSTM [25] CNN-LSTM [76] PEPNet [9] EventMamba
@@
TABLE VIII: Accuracy and Complexity on UCF101-DVS.
@@ 4.3.5 UCF101-DVS
For the larger 101-class dataset, EventMamba still performs strongly, reaching an impressive 97.1%, far exceeding all previous methods as shown in Table VIII. It is also very close to the 99.6% of the SOTA method [79] of UCF101 before synthesis. EVTC and Event-LSTM utilize the huge model and computational resources still only approach 80% accuracy. This is a great demonstration of the superiority of fine-grained temporal information and explicit temporal features. Additionally, we implement the dataset partitioning method based on filename, achieving accuracies of 60.4% and 90.3% for HMDB51-DVS and UCF101-DVS, respectively, still maintaining a leading performance level.
@@
In summary, our study employed the EventMamba model to conduct an all-encompassing evaluation of action recognition datasets, ranging from small to large scales. Our analysis demonstrated that EventMamba possesses a remarkable ability to generalize and achieve outstanding performance across datasets of varying sizes. However, after capturing explicit event temporal features, the model’s representational ability is greatly enhanced, and there is a greater risk of overfitting on small datasets, such as DVS Action and Daily DVS.
@@ 4.5 Results of Eye Tracking Dataset
Additionally, we conducted experiments with event-based eye tracking for the regression task. In contrast to the CPR task, which necessitates regression of translation and rotation in the world’s coordinate system, the eye tracking task solely requires regression of the pupil’s pixel coordinates within the camera coordinate system. To compare EventMamba’s performance more extensively, we use several representative Point Cloud networks [40, 11, 12, 9] for benchmarking. Table X illustrates that the optimal outcomes are highlighted in bold, whereas the second-best solutions are denoted in the underline. In comparison to classic Point Cloud networks, such as PointNet, PointNet++ and $\text{PointMLP}_{\text{elite}}$, EventMamba stands out by significantly enhancing accuracy and reducing distance error, and $\text{EventMamba}_{\text{lite}}$ also maintains high performance while consuming few computational resources. On the other hand, CNN, ConvLSTM, and CB-ConvLSTM eye-tracking frame-based methods [33] are selected for comparison. Our models exhibit a remarkable improvement in tracking accuracy, with a 3% increase in the $p_{3}$ and a 6% decrease in the MSE error, while maintaining a computational cost comparable to the aforementioned methods. Besides, compared to these frame-based methods, the parameters and FLOPs in EventMamba remain constant regardless of the input camera’s resolution. That enables EventMamba to be applied to different resolution edge devices.
4.6 Inference time
TABLE XIII: P5, P50, P95 inference time on diverse dataset.
Dataset P5 P50 P95
DVSGesture 5.6 ms 7.6 ms 9.5 ms
DVSAction 5.8 ms 6.0 ms 13.4 ms
DailyDVS 6.9 ms 7.2 ms 9.2 ms
HMDB51-DVS 7.1 ms 8.6 ms 15.0 ms
UCF101-DVS 7.3 ms 14.4 ms 15.7 ms
THU-CHL 7.1 ms 7.3 ms 8.4 ms
IJRR 5.0 ms 5.2 ms 6.2 ms
3ET 5.6 ms 5.9 ms 7.3 ms
We conducted tests on the P5, P50, and P95 inference time of EventMamba across all datasets on our server as shown in Table XIII. The server is equipped with a GTX 4090 GPU and an AMD 7950X CPU. This Table clearly shows that the inference time for action recognition is significantly shorter than the length of the sliding window, enabling high-speed real-time inference. Additionally, the throughput in eye tracking and camera pose relocation tasks far exceeds the 120 FPS frame rate of RGB cameras, ensuring that the high temporal resolution advantage of event cameras is maintained.
@@ 4.7.2 Point Number
In contrast to previous work [8, 58], we increased the number of input points to 2048 on large action recognition datasets, and to determine this significant impact factor, we verified it on HMDB51-DVS. The experiments compare the accuracy at training 100 epoch as described in Table XI, EventMamba with 2048 points has 82.1% accuracy on the testset, with 1024 points has 76.6% accuracy on the testset, and with 512 points has 67.9% accuracy on the testset. This shows that increasing the number of input points can significantly improve the performance of the model on large action recognition datasets.
@@ 4.7.3 Mamba or LSTM?
Table XI and XII show the performance of the two explicit feature extractors on the HMDB51-DVS and IJRR datasets, respectively. We try to ensure that both models consume an equitable computational resource, although the Mamba-based model still slightly surpasses the LSTM-based model. On the HMDB51-DVS dataset, Mamba demonstrated a notable performance advantage, achieving an accuracy of 82.1%, which exceeded LSTM’s accuracy of 70.8% by 16%. Likewise, on the IJRR dataset, Mamba exhibited superior performance with its average Mean Squared Error (MSE) outperforming LSTM’s by 8%. Additionally, the Mamba-based model’s inference is faster than LSTM, at 5.5 ms compared to 6.8 ms. Ignoring the slight increase in computational resources, we believe that Mamba is superior to LSTM in processing sequence events data.
@@ 5 Conclusion
In this paper, we rethink the temporal distinction between Event Cloud and Point Cloud and introduce a spatio-temporal hierarchical structure to address various event-based tasks. This innovative paradigm aligns with the inherently sparse and asynchronous characteristics of the event camera and effectively extracts implicit and explicit temporal information between events. Comprehensive experiments substantiate the efficacy of this approach in both regression and classification tasks, and we will derive this to more event-based tasks in the future.
6 Acknowledgment
This work was supported in part by the Young Scientists Fund of the National Natural Science Foundation of China (Grant 62305278), as well as the Guangdong Basic and Applied Basic Research Foundation (NO.2025A1515011758).